OpenStudy (anonymous):

solve for x: 2e^x=12

OpenStudy (chosenmatt):

one sec

OpenStudy (chosenmatt):

x=log(6)

OpenStudy (chosenmatt):

if you need anymore help just tag me :)

OpenStudy (chosenmatt):

ALSO WELCOME TO OPENSTUDY!!!

OpenStudy (anonymous):

log(6) has base e for further clarification; it is equivalent to ln(6)

OpenStudy (anonymous):

thank you so much! 2e^x=12 you divide the two which turns into e^x=6 which is log base e^6=x

OpenStudy (anonymous):

log base e ^6=x which is ln(6) right?

OpenStudy (chosenmatt):

YUP GOOD JOB!!

OpenStudy (anonymous):

thank you so much for your help!!

OpenStudy (chosenmatt):

if you need anything else just tag me :) and can you click best response please :)

OpenStudy (anonymous):

yeah no problem

OpenStudy (chosenmatt):

THX!!